# Greedy Algorithm for Structured Bandits: A Sharp Characterization of Asymptotic Success / Failure*

### v2

@@ Abstract
Version history. First version: March 2025. This version: May 2025 (adds Section 6 on infinite function classes).
@@ 1 Introduction
While exploration is central to bandit research, it presents significant challenges in practice, esp. when an algorithm interacts with human users. First, exploration can be wasteful and risky for the current user, imposing a burden that may be considered unfair since its benefits primarily accrue to future users. Second, exploration adds complexity to algorithm design,and its adoption in large-scale applications requires substantial buy-in and engineering support compared to a system that only exploits (Agarwal et al., 2016, 2017). Third, exploration may be incompatible with users’ incentives when actions are controlled by the users. E.g., an online platform cannot force users to try and review new products; instead, users gravitate toward well-reviewed or familiar options (Kremer et al., 2014). [^1] While exploration can be made incentive-compatible in such settings, doing so introduces additional cost and complexity (Kremer et al. (2014) and follow-up work, see Slivkins (2023) for an overview).
@@
Despite its simplicity and practical appeal, $\mathtt{Greedy}$ is widely believed to perform poorly. This belief is deeply ingrained in the bandit literature, which overwhelmingly focuses on exploration as a necessary ingredient for minimizing regret. A key motivation for this focus comes from well-known failure cases in unstructured $K$-armed bandits. A classic example is as follows: “Suppose the reward of each arm follows an independent Bernoulli distribution with a fixed mean, and $\mathtt{Greedy}$ is initialized with a single sample per arm. If the best arm initially returns a 00 while another arm returns a $1$, $\mathtt{Greedy}$ permanently excludes the best arm.”
However, beyond such examples, the broader picture remains murky, especially for the widely-studied structured bandits—bandit problems with a known reward structure (e.g., linearity, Lipschitzness, convexity)—where observing some actions provides useful information about others. Formally, a reward structure restricts the possible reward functions that map arms to their mean rewards. Reward structures reduce the need for explicit exploration, making the bandit problem more tractable. For some of them, $\mathtt{Greedy}$ in fact succeeds, e.g., two-armed bandits with expected rewards that sum up to a known value. The literature provides a few examples of failure for some specific (one-dimensional, linear) reward structures, and a few non-trivial examples of success (e.g., for linear contextual bandits), more on this in Related Work. Likewise, large-scale experiments yield mixed results: some settings confirm the need for exploration, but others indicate that $\mathtt{Greedy}$ performs well even (Bietti et al., 2021). This contrast raises a fundamental question: When—and why—does $\mathtt{Greedy}$ fail or succeed?
@@
A key insight is identifying a new “partial identifiability” property of the problem instance, called self-identifiability, as a necessary and sufficient condition for the asymptotic success. Self-identifiability asserts that, given the reward structure, fixing the expected reward of a suboptimal arm uniquely identifies it as suboptimal. We prove that $\mathtt{Greedy}$ achieves sublinear regret under self-identifiability, and suffers from linear regret otherwise. The negative result is driven by the existence of a decoy: informally, an alternative reward model such that its optimal arm is suboptimal for the true model and both models coincide on this arm. We show that with some positive probability, $\mathtt{Greedy}$ gets permanently stuck on such decoy, for an infinite time horizon.
$\mathtt{Greedy}$ succeeds (only) because self-identifiability makes the problem instance intrinsically “easy”: any reasonable algorithm (under a mild non-degeneracy condition) achieves sublinear regret.
@@
Our characterization allows for essentially an arbitrary interaction protocol between the algorithm and the environment (Section 5). Specifically, we handle interactive decision-making with arbitrary auxiliary feedback (Foster et al., 2021), which subsumes contextual bandits and combinatorial semi-bandits, as well as episodic reinforcement learning. That said, before moving to this full generality, our presentation focuses on contextual bandits, where we obtain quantitatively stronger guarantees (Section 4), and “vanilla” bandits as a paradigmatic case for building key intuition (Section 3).
The second main result of this work is a partial characterization which handles arbitrary infinite (e.g., continuous) reward structures. It applies to structured bandits with finite action sets, requiring stronger notions of self-identifiability and decoy existence, as well as new analysis ideas, see Section 6.
@@
We apply our machinery to several examples, both positive and negative (Section 7). We find that $\mathtt{Greedy}$ fails in linear bandits, Lipschitz bandits and “polynomial bandits” (with arms in $\mathbb{R}$ and polynomial expected rewards), and does so for almost all problem instances. For linear contextual bandits, $\mathtt{Greedy}$ succeeds if the context set is “sufficiently diverse”, but may fail if it is “low-dimensional”. For Lipschitz contextual bandits, $\mathtt{Greedy}$ behaves very differently, failing for almost all instances. One informal takeaway is that $\mathtt{Greedy}$ fails as a common case for most/all bandit structures of interest, whereas for contextual bandits it can go either way. The success of $\mathtt{Greedy}$ appears to require context diversity and a parametric reward structure.
Discussion. While our first main result—the complete characterization—assumes finite reward structures, most infinite structures of interest admit meaningful finite analogs via discretization (see Section 2 and illustrations in Section 7). Moreover, we relax the finiteness assumption in Section 6.
@@
The linear vs sublinear regret is a standard “first-order” notion of success/failure in bandits. Our positive results attain logarithmic worst-case regret rates, possibly with a large multiplicative constant determined by the reward structure. Our negative results establish a positive-constant (but possibly very small) probability of a “failure event” when $\mathtt{Greedy}$ gets permanently stuck on a decoy, for an infinite time horizon. Optimizing these constants for an arbitrary reward structure appears difficult. However, we achieve much better constants for the partial characterization in Section 6.
@@
For $\mathtt{Greedy}$, positive results with near-optimal regret rates focus on linear contextual bandits with diverse/smoothed contexts (Kannan et al., 2018; Bastani et al., 2021; Raghavan et al., 2023; Kim and Oh, 2024). The greedy algorithm is also known to attain $o(T)$ regret in various scenarios with a very large number of near-optimal arms (Bayati et al., 2020; Jedor et al., 2021). [^2] E.g., for Bayesian bandits with $\gg\sqrt{T}$ arms, where the arms’ mean rewards are sampled uniformly.
Negative results for $\mathtt{Greedy}$ are derived for “non-structured” $K$-armed bandits: from trivial extensions of the single-sample-per-arm example mentioned above, to an exponentially stronger characterization of failure probability (Banihashem et al., 2023a), to various “near-greedy” algorithms / behaviors, both “frequentist” and “Bayesian” (same paper). Negative results for non-trivial reward structures concern dynamic pricing with linear demands (Harrison et al., 2012; den Boer and Zwart, 2014) and dynamic control in a generalized linear model (Lai and Robbins, 1982; Keskin and Zeevi, 2018). Banihashem et al. (2023a, b) also obtain negative results for the Bayesian version of $\mathtt{Greedy}$ in Bayesian bandits, under a certain “full support” assumption on the prior. [^3] Essentially, the prior allows all reward functions $\{\text{arms}\}\to[0,1]$ with probability density at least $p>0$.
@@ 2 Preliminaries: structured contextual bandits ($\mathtt{StructuredCB}$)
We have action set $\mathcal{A}$ and context set $\mathcal{X}$. In each round $t=1,2,\,\ldots$, a context $x_{t}\in\mathcal{X}$ arrives, an algorithm chooses an action (arm) $a_{t}\in\mathcal{A}$, and a reward $r_{t}\in\mathbb{R}$ is realized. The context is drawn independently from some fixed and known distribution over $\mathcal{X}$.  [^4] Whether the context distribution is known to the algorithm is inconsequential, since $\mathtt{Greedy}$ (particularly, the regression in Eq. (2.1)) does not use this knowledge. W.l.o.g., $\mathcal{X}$ is the support set of the context distribution. The reward $r_{t}$ is an independent draw from a unit-variance Gaussian with unknown mean $f^{*}(x_{t},a_{t})\in[0,1]$.  [^5] Gaussian reward noise is a standard assumption in bandits (along with e.g., 0-1 rewards), which we make for ease of presentation. Our positive results carry over to rewards with an arbitrary sub-Gaussian noise, without any modifications. Likewise, our negative results carry over to rewards $r_{t}\in[0,1]$ with an arbitrary near-uniform distribution, i.e., one specified by a p.d.f. on $[0,1]$ which is bounded away from 00 by an absolute constant. A reward function is a function $f:\mathcal{X}\times\mathcal{A}\to[0,1]$; in particular, $f^{*}$ is the true reward function. The reward structure is given by a known class $\mathcal{F}$ of reward functions which contains $f^{*}$; the assumption $f^{*}\in\mathcal{F}$ is known as realizability. To recap, the problem instance is a pair $(f^{*},\mathcal{F})$, where $\mathcal{F}$ is known and $f^{*}$ is not.
@@ 3 Characterization for structured bandits
$\Gamma(f^{*},\mathcal{F})=\min_{\textrm{functions }f\in\mathcal{F}:\,f\neq f^{*}}\quad\min_{\textrm{arms }a:\,f(a)\neq f^{*}(a)}\quad\absolutevalue{f^{*}(a)-f(a)}.$ | . (3.1)
@@
$E_{1}=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(a)-f_{\mathtt{dec}}(a)}<\Gamma(f_{\mathtt{dec}})/2\quad\text{for each arm $a\neq a_{\mathtt{dec}}$}\,\right\}.$ (3.2)
@@
$E_{2}=\left\{\,\forall t>T_{0},\;\absolutevalue{\bar{r}_{t}(a_{\mathtt{dec}})-f^{*}(a_{\mathtt{dec}})}\leq\Gamma(f_{\mathtt{dec}})/2\,\right\}.$ (3.3)
@@ 4 Characterization for structured contextual bandits ($\mathtt{StructuredCB}$)
$\Gamma(f^{*},\mathcal{F})=\min_{\textrm{functions }f\in\mathcal{F}:\,f\neq f^{*}}\quad\min_{(x,a)\in\mathcal{X}\times\mathcal{A}:\;f(x,a)\neq f^{*}(x,a)}\quad\absolutevalue{f^{*}(x,a)-f(x,a)}.$ | . (4.1)
@@ 5 Interactive decision-making with arbitrary feedback
We consider Decision-Making with Structured Observations ($\mathtt{DMSO}$), a general framework for sequential decision-making with a known structure (Foster et al., 2021). It allows for arbitrary feedback observed after each round, along with the reward. [^7] Bandit formulations with partial feedback that does not include the reward (known as partial monitoring, e.g.,  Bartók et al., 2014; Antos et al., 2013), are outside our scope. This feedback may be correlated with the resp. rewards, necessitating a modification of $\mathtt{Greedy}$ which looks beyond average/expected rewards. The analysis becomes considerably more technical compared to $\mathtt{StructuredCB}$.
@@
Preliminaries. $\mathtt{DMSO}$ is defined as follows. Instead of “arms” and “contexts”, we have two new primitives: a policy set $\Pi$ and observation set $\mathcal{O}$. The interaction protocol is as follows: in each round $t=1,2,\,\ldots$, the algorithm selects a policy $\pi_{t}\in\Pi$, receives a reward $r_{t}\in\mathcal{R}\subset\mathbb{R}$, and observes an observation $o_{t}\in\mathcal{O}$. A model is a mapping from $\Pi$ to a distribution over ×$\mathbb{R}\times\mathcal{O}$. The reward-observation pair $(r_{t},o_{t})$ is an independent sample from distribution $M^{*}(\pi_{t})$, where $M^{*}$ is the true model. The problem structure is represented as a (known) model class $\mathcal{M}$ which contains $M^{*}$. We assume that $\Pi,\mathcal{M},\mathcal{R},\mathcal{O}$ are all finite. [^8] Finiteness of $\mathcal{R},\mathcal{O}$ is for ease of presentation. We can also handle infinite $\mathcal{R},\mathcal{O}$ if all outcome distributions $M(\pi)$ have a well-defined density, and 1 is stated in terms of these densities. To recap, the problem instance is a pair $(M^{*},\mathcal{M})$, where $\mathcal{M}$ is known but $M^{*}$ is not. This completes the definition of $\mathtt{DMSO}$.
@@
$\mathtt{StructuredMAB}$ is a simple special case of $\mathtt{DMSO}$ with one possible observation. $\mathtt{StructuredCB}$ is subsumed by interpreting the observations $o_{t}$ as contexts and defining $M^{*}(\pi)$ accordingly, to account for the distribution of context arrivals, the reward distribution, and the reward function. [^9] Here we work with discrete rewards, whereas our treatment in Sections 3 and 4 assumes Gaussian rewards. The observations in $\mathtt{DMSO}$ can also include auxiliary feedback such as, e.g., rewards of “atomic actions” in combinatorial semi-bandits (e.g.,  György et al., 2007; Chen et al., 2013), per-product sales in multi-product dynamic pricing (e.g.,  Keskin and Zeevi, 2014; den Boer, 2014), and MDP trajectories in episodic reinforcement learning (see Agarwal et al. (2020) for background). $\mathtt{DMSO}$ subsumes all these scenarios, under the “realizability” assumption $M^{*}\in\mathcal{M}$.
@@
happens w.h.p. when $\mathcal{H}_{\mathtt{warm}}$ comes from $\mathtt{ghost}$ [^13] This argument invokes a concentration inequality, which in turn uses 1. Likewise, 1 is used for another application of concentration in the end of the proof sketch. Since $\mathtt{ghost}$ and $\mathcal{H}_{\mathtt{warm}}$ have bounded Renyi divergence, we argue that with some positive probability, event $E_{1}$ happens under $\mathcal{H}_{\mathtt{warm}}$.

### v3

@@ Abstract
Version history. 2025/03: first version. 2025/05: Section 6 on infinite function classes. 2025/11: revised presentation.
@@ 1 Introduction
While exploration is central to bandit research, it presents significant challenges in practice, esp. when an algorithm interacts with human users. First, exploration can be wasteful and risky for the current user, imposing a burden that may be considered unfair since its benefits primarily accrue to future users. Second, exploration adds complexity to algorithm design,and its adoption in large-scale applications requires substantial buy-in and engineering support compared to a system that only exploits (Agarwal et al., 2016, 2017). Third, exploration may be incompatible with users’ incentives when actions are controlled by the users. E.g., an online platform cannot force users to try and review new products; instead, users gravitate toward well-reviewed or familiar options (Kremer et al., 2014). [^1] Enforcing exploration in such settings is very challenging (Kremer et al. (2014) and follow-up work, see Slivkins (2023) for an overview). While exploration can be made incentive-compatible, doing so involves considerable performance and/or monetary costs and additional complexity. More importantly, it hinges upon substantial (even if standard) assumptions from economic theory.
@@
Despite its simplicity and practical appeal, $\mathtt{Greedy}$ is widely believed to perform poorly. This belief is deeply ingrained in the bandit literature, which overwhelmingly focuses on exploration as a necessary ingredient for minimizing regret. A key motivation for this focus comes from well-known failure cases in unstructured $K$-armed bandits. A classic example is as follows: “Suppose the reward of each arm follows an independent Bernoulli distribution with a fixed mean, and $\mathtt{Greedy}$ is initialized with a single sample per arm. If the best arm initially returns a 0 while another arm returns a $1$, $\mathtt{Greedy}$ permanently excludes the best arm.”
However, beyond such examples, the broader picture remains murky, especially for the widely-studied structured bandits – bandit problems with a known reward structure (e.g., linearity, Lipschitzness, convexity) – where observing some actions provides useful information about others. Formally, a reward structure restricts the possible reward functions that map arms to their mean rewards. Reward structures reduce the need for explicit exploration, making the bandit problem more tractable. For some of them, $\mathtt{Greedy}$ in fact succeeds, e.g., two-armed bandits with expected rewards that sum up to a known value. The literature provides a few examples of failure for some specific (one-dimensional, linear) reward structures, and a few non-trivial examples of success (e.g., for linear contextual bandits); see more on this in Related Work. Likewise, large-scale experiments yield mixed results: some settings confirm the need for exploration, but others indicate that $\mathtt{Greedy}$ performs well (Bietti et al., 2021). This contrast raises a fundamental question: When—and why—does $\mathtt{Greedy}$ fail or succeed?
@@
A key insight is identifying a new “partial identifiability” property of the problem instance, called self-identifiability, as a necessary and sufficient condition for the asymptotic success. Self-identifiability asserts that, given the reward structure, fixing the expected reward of a suboptimal arm uniquely identifies it as suboptimal. We prove that $\mathtt{Greedy}$ achieves sublinear regret under self-identifiability, and suffers from linear regret otherwise. The negative result is driven by the existence of a decoy: informally, an alternative reward model such that its optimal arm is suboptimal for the true model and both models coincide on this arm. We show that with some positive probability, $\mathtt{Greedy}$ gets permanently stuck on such a decoy, for an infinite time horizon. For the positive result, $\mathtt{Greedy}$ succeeds (only) because self-identifiability makes the problem instance intrinsically “easy”. In fact, this success is not due to any particular cleverness of $\mathtt{Greedy}$: we show that any reasonable algorithm (satisfying a mild non-degeneracy condition) achieves sublinear regret under self-identifiability.
@@
Our characterization allows for essentially an arbitrary interaction protocol between the algorithm and the environment (Section 5). Specifically, we handle the model of “decision-making with structured observations” ($\mathtt{DMSO}$, Foster et al. (2021)), which allows for arbitrary auxiliary feedback after each round. This model subsumes contextual bandits, combinatorial semi-bandits, and bandits with graph-based feedback, as well as episodic reinforcement learning. Before moving to this full generality, our presentation focuses on contextual bandits, where we obtain quantitatively stronger guarantees (Section 4), and “vanilla” bandits as a paradigmatic case for building key intuition and cleaner definitions (Section 3).
@@
We apply our machinery to several examples, both positive and negative (Section 7). We demonstrate that most infinite structures of interest admit meaningful finite analogs via discretization. We find that $\mathtt{Greedy}$ fails in linear bandits, Lipschitz bandits and “polynomial bandits” (with arms in $\mathbb{R}$ and polynomial expected rewards), and does so for almost all problem instances. For linear contextual bandits, $\mathtt{Greedy}$ succeeds if the context set is “sufficiently diverse”, but may fail if it is “low-dimensional”. For Lipschitz contextual bandits, $\mathtt{Greedy}$ behaves very differently, failing for almost all instances. One informal takeaway is that $\mathtt{Greedy}$ fails as a common case for most/all bandit structures of interest, whereas for contextual bandits it can go either way, depending on the structure. The success of $\mathtt{Greedy}$ appears to require context diversity and a parametric reward structure.
The second main result of this work concerns infinite (e.g., continuous) reward structures (Section 6). While our earlier analysis gave a sharp “if and only if” characterization for finite reward structures, such a complete characterization is more challenging to obtain for infinite ones. To make progress, we provide a characterization parameterized by a notion of “margin” (separating instances for which the positive result applies from instances for which the negative result applies), with guarantees that deteriorate as the margin vanishes. Subject to this margin, our result handles arbitrary infinite reward structures. It applies to structured bandits with finite action sets, requiring stronger notions of self-identifiability and decoy existence (parameterized by the margin), as well as new analysis ideas.
@@
Discussion. The distinction between linear and sublinear regret is a standard “first-order” notion of success vs failure in bandits. Our positive results attain logarithmic, instance-dependent regret rates, possibly with a large multiplicative constant determined by the reward structure and the instance. Our negative results establish a positive (but possibly very small) constant probability of a “failure event” where $\mathtt{Greedy}$ gets permanently stuck on a decoy, for an infinite time horizon. Optimizing these constants for an arbitrary reward structure appears difficult. However, we achieve much better constants for the partial characterization in Section 6.
@@
For $\mathtt{Greedy}$, positive results with near-optimal regret rates focus on linear contextual bandits with diverse/smoothed contexts (Kannan et al., 2018; Bastani et al., 2021; Raghavan et al., 2023; Kim and Oh, 2024). Both context diversity and parametric reward structure are essential. Our positive results for the same setting are incomparable: more general in terms of context diversity assumptions, but weaker in terms of the regret bounds. (This line of prior work does not contain any negative results.) $\mathtt{Greedy}$ is also known to attain $o(T)$ regret in various scenarios with a very large number of near-optimal arms (Bayati et al., 2020; Jedor et al., 2021). [^2] E.g., for Bayesian bandits with $\gg\sqrt{T}$ arms, where the arms’ mean rewards are sampled uniformly.
Negative results for $\mathtt{Greedy}$ are derived for “non-structured” $K$-armed bandits: from trivial extensions of the single-sample-per-arm example mentioned above, to an exponentially stronger characterization of failure probability (Banihashem et al., 2023), to various “near-greedy” algorithms / behaviors, both “frequentist” and “Bayesian” (same paper). Negative results for non-trivial reward structures concern dynamic pricing with linear demands (Harrison et al., 2012; den Boer and Zwart, 2014) and dynamic control in a generalized linear model (Lai and Robbins, 1982; Keskin and Zeevi, 2018). Banihashem et al. (2023) also obtain negative results for the Bayesian version of $\mathtt{Greedy}$ in Bayesian bandits, under a certain “full support” assumption on the prior. [^3] Essentially, the prior covers all reward functions $\{\text{arms}\}\to[0,1]$ with probability density at least $p>0$.
@@ 2 Preliminaries: structured contextual bandits ($\mathtt{StructuredCB}$)
We have action set $\mathcal{A}$ and context set $\mathcal{X}$. In each round $t=1,2,\,\ldots$, a context $x_{t}\in\mathcal{X}$ arrives, an algorithm chooses an action (arm) $a_{t}\in\mathcal{A}$, and a reward $r_{t}\in\mathbb{R}$ is realized. The context is drawn independently from some fixed and known distribution over $\mathcal{X}$.  [^4] Whether the context distribution is known to the algorithm is inconsequential, since $\mathtt{Greedy}$ (particularly, the regression in Eq. (2.1)) does not use this knowledge. W.l.o.g., $\mathcal{X}$ is the support set of the context distribution. The reward $r_{t}$ is an independent draw from a unit-variance Gaussian with unknown mean $f^{*}(x_{t},a_{t})\in[0,1]$.  [^5] Gaussian reward noise is a standard assumption in bandits (along with e.g., 0-1 rewards), which we make for ease of presentation. Our positive results carry over to rewards with an arbitrary sub-Gaussian noise, without any modifications. Likewise, our negative results carry over to rewards $r_{t}\in[0,1]$ with an arbitrary near-uniform distribution, i.e., one specified by a p.d.f. on $[0,1]$ which is bounded away from 0 by an absolute constant. A reward function is a function $f:\mathcal{X}\times\mathcal{A}\to[0,1]$; in particular, $f^{*}$ is the true reward function. The reward structure is given by a known class $\mathcal{F}$ of reward functions which contains $f^{*}$; the assumption $f^{*}\in\mathcal{F}$ is known as realizability. To recap, the problem instance is a pair $(f^{*},\mathcal{F})$, where $\mathcal{F}$ is known and $f^{*}$ is not.
@@ 3 Characterization for structured bandits
$|.\displaystyle\Gamma(f^{*},\mathcal{F})=\min_{\textrm{functions }f\in\mathcal{F}:\,f\neq f^{*}}\quad\min_{\textrm{arms }a:\,f(a)\neq f^{*}(a)}\quad\absolutevalue{f^{*}(a)-f(a)}$. (3.1)
@@
[…]
@@
[…]
@@ 4 Characterization for structured contextual bandits ($\mathtt{StructuredCB}$)
$|.\displaystyle\Gamma(f^{*},\mathcal{F})=\min_{\textrm{functions }f\in\mathcal{F}:\,f\neq f^{*}}\quad\min_{(x,a)\in\mathcal{X}\times\mathcal{A}:\;f(x,a)\neq f^{*}(x,a)}\quad\absolutevalue{f^{*}(x,a)-f(x,a)}$. (4.1)
@@ 5 Interactive decision-making with arbitrary feedback
We consider Decision-Making with Structured Observations ($\mathtt{DMSO}$), a general framework for sequential decision-making with a known structure (Foster et al., 2021). It allows for arbitrary feedback observed after each round, along with the reward. [^7] Bandit formulations with partial feedback that does not include the reward (known as partial monitoring, e.g.,  Bartók et al., 2014; Antos et al., 2013), are outside our scope. The root challenge is that this feedback is usually correlated with rewards. $\mathtt{Greedy}$ must account for these correlations, not just compute the best fit based on rewards. Our solution is to develop a natural variant of $\mathtt{Greedy}$ based on maximum-likelihood estimation. The analysis then becomes much more technical compared to $\mathtt{StructuredCB}$, requiring us to track changes in log-likelihood and define the ”model-gap” in terms of KL-divergence.
@@
Preliminaries. $\mathtt{DMSO}$ is defined as follows. Instead of “arms” and “contexts”, we have two new primitives: a policy set $\Pi$ and observation set $\mathcal{O}$. The interaction protocol is as follows: in each round $t=1,2,\,\ldots$, the algorithm selects a policy $\pi_{t}\in\Pi$, receives a reward $r_{t}\in\mathcal{R}\subset\mathbb{R}$, and observes an observation $o_{t}\in\mathcal{O}$. A model is a mapping from $\Pi$ to a distribution over $\mathbb{R}\times\mathcal{O}$. The reward-observation pair $(r_{t},o_{t})$ is an independent sample from distribution $M^{*}(\pi_{t})$, where $M^{*}$ is the true model. The problem structure is represented as a (known) model class $\mathcal{M}$ which contains $M^{*}$. We assume that $\Pi,\mathcal{M},\mathcal{R},\mathcal{O}$ are all finite. [^8] Finiteness of $\mathcal{R},\mathcal{O}$ is for ease of presentation. We can also handle infinite $\mathcal{R},\mathcal{O}$ if all outcome distributions $M(\pi)$ have a well-defined density, and 1 is stated in terms of these densities. To recap, the problem instance is a pair $(M^{*},\mathcal{M})$, where $\mathcal{M}$ is known but $M^{*}$ is not. This completes the definition of $\mathtt{DMSO}$.
@@
$\mathtt{StructuredMAB}$ is a simple special case of $\mathtt{DMSO}$ with one possible observation. $\mathtt{StructuredCB}$ is subsumed by interpreting the observations $o_{t}$ as contexts and defining $M^{*}(\pi)$ accordingly, to account for the distribution of context arrivals, the reward distribution, and the reward function. [^9] Here we work with discrete rewards, whereas our treatment in Sections 3 and 4 assumes Gaussian rewards. The observations in $\mathtt{DMSO}$ can also include auxiliary feedback present in various bandit models studied in prior work. To wit: rewards of “atomic actions” in combinatorial semi-bandits (e.g.,  György et al., 2007; Chen et al., 2013), per-product sales in multi-product dynamic pricing (e.g.,  Keskin and Zeevi, 2014; den Boer, 2014), and rewards of all “adjacent” arms in bandits with graph-based feedback (Alon et al., 2013, 2015). Moreover, the observations can include MDP trajectories in episodic reinforcement learning (see Agarwal et al., 2020, for background). $\mathtt{DMSO}$ subsumes all these scenarios, under the “realizability” assumption $M^{*}\in\mathcal{M}$.
@@
happens w.h.p. when $\mathcal{H}_{\mathtt{warm}}$ comes from $\mathtt{ghost}$. [^13] This argument invokes a concentration inequality, which in turn uses 1. Likewise, 1 is used for another application of concentration in the end of the proof sketch. Since $\mathtt{ghost}$ and $\mathcal{H}_{\mathtt{warm}}$ have bounded Renyi divergence, we argue that with some positive probability, event $E_{1}$ happens under $\mathcal{H}_{\mathtt{warm}}$.